OpenStudy (anonymous):

Choose the equation below that represents the line that passes through the point (-2, -1) and has a slope of 5. y - 1 = 5(x - 2) y + 1 = 5(x + 2) y + 2 = 5(x + 1) y - 2 = 5(x - 1)

OpenStudy (anonymous):

The equation with one point given and given slope is given by: \[\huge y - y_1 = m(x-x_1)\] Here m= 5 and \((x_1, y_1)\) = (-2, -1) Put and find the equation..

OpenStudy (anonymous):

Is it D.?

OpenStudy (anonymous):

See, -(-) becomes +.. \[\large y- (-1) = 5(x - (-2))\] Now solve it..

OpenStudy (anonymous):

Its A. or B.

OpenStudy (anonymous):

Is it A

OpenStudy (anonymous):

It is B..

OpenStudy (anonymous):

-(-) becomes + So, y + 1 = 5(x + 2)